# Probing the axial symmetry of gamma-ray burst jets using afterglow polarimetry

### v1

@@ 1 Introduction
Numerical simulations suggest such jet structures can arise from the interaction between the jet and the confining medium (Gottlieb et al., 2021). Specifically, this occurs as the jet breaks out through the stellar envelope in the collapsar model (e.g. Zhang et al. 2003; Lazzati & Begelman 2005; Gottlieb et al. 2021); or interacts with merger ejecta (e.g. Perego et al. 2014; Nativi et al. 2022). While most afterglow studies to date have focused on axisymmetric jets (either top-hat or structured), some works have explored with uneven energy and velocity distributions in the azimuthal direction (Li et al. 2023; Li et al. 2024), which discuss uneven distribution of energy and velocity within a jet in the azimuth angle direction due to internal non-uniform magnetic dissipation processes or the precession of the central engine (Huang et al., 2019). Recent three-dimensional hydrodynamic simulations of jets in the aftermath of neutron star mergers (Gottlieb et al. 2022; Lamb et al. 2022) further highlight these complexities. These studies show that when a jet breaks out of the ejecta, the resulting outflow exhibits a power-law-like polar energy distribution with rotational inhomogeneity. Notably, the jet head is not perfectly circular, as often assumed in conventional theoretical models, instead, it tends to exhibit a more elliptical shape.
Non-axisymmetric or inhomogeneous structures have been primarily discussed in the context of prompt gamma-ray emission (e.g. Gill & Granot 2024; Lazar et al. 2009). This emission originates from the original ejecta from the central engine, which carries large-scale magnetic fields generated at the central source. While polarisation signals in the prompt gamma-rays are more sensitive to the magnetic field structure within the jet, those in the afterglow phase are predominantly influenced by geometrical effects. Consequently, afterglow polarimetry offers a more effective means of probing the jet’s structure and geometry.
@@
We aim to investigate the signature of jet axial asymmetry imprinted on the temporal evolution of afterglow polarisation. To break the axisymmetry, we model the global jet structure with an elliptical head. While real jets may exhibit more complex geometries, this approach represents a natural starting point for exploring non-axisymmetric jets in this context. We present our model for evaluating polarisation signals around a jet break in Section 2, where we illustrate the primary polarisation signature of a non-axisymmetric jet. We describe the distributions of polarisation signals when the line-of-sight of an observer intersects the jet head (or shock surface) at random locations in Section 3. In Section 4, we discuss the structured model in the context of non-axisymmetric jets. In Section 5 we present case studies. Finally, in Section 6, we give our conclusions and discussion.
@@ 2 Polarisation Behaviour at a Jet Break
Figure 2: Top panel: Polarisation as a function of time for $e=0$ (blue dashed) and 0.6 (red solid). The polarisation scales almost linearly with the local polarisation degree $P_{0}$, and we have assumed $P_{0}=0.7$. The first peaks, the troughs and second peaks in the polarisation light curves are marked by circles, crosses and squares, respectively. The line of sight is characterised by $\chi=1$ and $\xi=0.3$. Bottom panel: Same as the top panel, but showing the the evolution of polarisation signals in the Stokes parameter space $(q,u)$.
In the afterglow phase, magnetic fields in the blast wave are believed to be produced by shock instabilities, and expected to be highly tangled. However, the precise mechanisms by which magnetic fields are produced in relativistic collisionless shocks remain poorly understood. Since the shock normal (aligned with the radial direction) represents a distinct orientation, the magnetic field components parallel and perpendicular to the shock normal could exhibit significantly different average strengths (Medvedev & Loeb 1999; Sari 1999; Ghisellini & Lazzati 1999).
@@
where the local polarisation degree $P_{e}=P_{0}$ when the segment lies within the ring, and $P_{e}=0$ when the segment is inside the inner circle. The local luminosity, $dL$, is proportional to the area of the segment while also incorporating the constraints imposed by conditions (3), (5), and (6) in the previous paragraph. The angle $\phi_{e}$ denotes the angle between the positive x-axis (the semi-major axis) and the segment’s position vector, measured relative to the light of sight (Point A in Figure 1). Considering condition (6), $\Sigma dQ$, $\Sigma dU$ and $\Sigma dL$ are evaluated to estimate the net values of the parameters: $q=\Sigma dQ/\Sigma dL$ and $u=\Sigma dU/\Sigma dL$. The net polarisation degree $P$ and position angle $\phi$ are given by
@@
Figure 2 illustrates the differences in polarisation signals between axisymmetric jets and non-axisymmetric jets (in this paper, non-axisymmetric jets refer to those with elliptical-shaped heads). For axisymmetric jets, the polarisation degree curve (shown as the blue dashed line in the top panel) exhibits two peaks, with the polarisation dropping to zero between the peaks around the jet break time. As seen in the bottom panel, the time evolution of $(q,u)\propto(\cos 2\phi,\sin 2\phi)$ follows a linear trajectory passing through the origin. The polarisation position angle (PPA) $\phi$ rotates precisely by $\pi/2$ before and after the point where polarisation vanishes, around the jet break time (Sari 1999; Ghisellini & Lazzati 1999). In contrast, jets with elliptical heads generally do not exhibit a complete polarisation drop between the two peaks ($P\neq 0$ at the trough). The PPA rotates more smoothly, though it still undergoes rapid changes near the trough (approximately at the jet break time). The change in PPA between the two peaks can significantly deviate from $\pi/2$; in this example, it is approximately $\Delta\phi=0.28\pi\sim 50.4$ degrees (it is more convenient to express angles in radians for theoretical discussion, while degrees are preferred in the context of observations. Unless otherwise specified, angles are given in radians).
@@ 3 The Characterisation of polarisation signals
Similarly, consider Point C, which is symmetric to Point A with respect to the x-axis. The fluid element distribution around Point C mirrors that of Point A but with a reflection of $\phi\leftrightarrow-\phi$ (flipping vertically with respect to the line passing through Point A and parallel to the x-axis). While the absolute values of the Stokes parameters at Point C match those at Point A, one of them flips sign, specifically $(q,-u)$. Nevertheless, the position angle rotation $|\Delta\phi-\pi/2|$ and the ratio $R_{pd}$ remain the same as at Point A.
The jet head is a part of a spherical blast wave, but the ellipse represented in Figure 2 represents its projection onto the sky. The opening angle of GRB jets or their core size is typically only several degrees. For such small angles, uniform sampling over projected area is approximately equivalent to uniform sampling over the solid angle of the shock surface.
Figure 3 shows the distributions of $|\Delta\phi-\pi/2|$, derived from 5000 uniform sampled line-of-sight realisations across a quadrant of the elliptic jet head for eccentricity $e=0.3$, 0.45, and 0.6. For axisymmetric jets ($e=0$), the probability distribution is given by a delta function at 0, meaning the PPA rotation $\Delta\phi$ is always $\pi/2$. As eccentricity increases, the probability distribution (solid lines) peaks at larger deviation angles, with the distribution’s spread also increasing. The probability density distributions peak at $|\Delta\phi-\pi/2|$ = 4.3, 9.5 and 19 degrees, while their cumulative distributions reach 50% at 3.8, 8.9, 18 degrees, and 90% at 7.3, 17, 31 degrees for $e=0.3,0.45$ and 0.6, respectively. Since the semi-minor axis is smaller than the semi-major axis by a factor of $\sqrt{1-e^{2}}$, it is reduced by approximately 5%–20% (or equivalently by a factor of $\sim 0.8-0.95$) across this range of eccentricities. Our results show that even a mild deformation of the jet head leads to PPA rotations that deviate significantly from $\pi/2$.
@@
For even smaller offsets, the first peak becomes comparable to the trough, resulting in $R_{pd}\sim 1$. To determine whether a polarisation curve exhibits two peaks and to locate them, we have implemented a peak-finding algorithm similar to the one discussed in Li & Fenimore 1996 (see Appendix A for more details). For approximately, 4%, 7% and 11% of uniform random realisations for the $e=0.3$, 0.45 and 0.6, respectively, the polarisation curves exhibit only a single peak. These single peak cases have been excluded from the analysis presented in Figs 3, 4 and 5.
@@ 4 Non-axisymmetric structured jet
Figure 6: Polarisation signals for a line of sight intersecting the outer region, where the surface luminosity density declines following a power-law with an index of $\alpha=2$ (red) or 3 (blue). Two offset angles are considered: $\xi=2\theta_{c}/\theta_{j}$ (solid) and $4\theta_{c}/\theta_{j}$ (dashed). Top panel: temporal evolution of polarisation degree. Bottom panel: the evolution of the Stokes parameter $(q,u)$. The model assumes $e=0.3$, $\chi=\pi/4$, and $\theta_{c}/\theta_{j}=0.13$.
Figure 7: Polarisation signals for a line of sight intersecting the jet core at $\xi=0.5\theta_{c}/\theta_{j}$. All jet properties are identical to those in Figure 6, except where specified. Top panel: polarisation degree curves. Bottom panel: evolution of the Stoke parameters. Curves are shown for $\alpha=2$ (red) and $\alpha=3$ (blue). The first peaks, the troughs and second peaks in the polarisation light curves are marked by circles, crosses and squares, respectively.
@@
Figure 6 shows the polarisation signals when the line of sight falls in the outer region. As demonstrated by Rossi et al. 2004 and Birenbaum et al. 2024, the polarisation curves differs significantly from those of a top-hat jet, regardless of axial symmetry. This discrepancy arises from the non-uniform luminosity distribution in the visible region surrounding the line of sight (i.e., the ring), with the intense emission concentrated toward the jet axis. Consequently, the observed polarisation signal is primarily influenced by these segments of the ring, resulting in a single peak in the polarisation curve, which occurs around the time the core becomes visible to the observer (top panel). The polarisation degree curves for the decay indices $\alpha=2$ and 3 exhibit similar shapes.
@@
Since the the emission from the ring segments closer to the jet axis dominates the polarisation signals, the position angle evolution also differs from those of a top-hat jet. It consistently aligns with the direction toward the central jet axis (see assumption (4) in Section 2). When the line of sight falls on the jet head near its semi-major or semi-minor axis ($\chi\sim 0$ or $\pi/2$), the geometry becomes more symmetric, the position angle remains constant over time, consistent with the results for axisymmetric structured jets (Rossi et al., 2004). However, for a general azimuthal angle $\chi$, the eccentricity of the brightness contour lines leads to a slight evolution of the position angle around the peak of the polarisation degree. This effect is illustrated for $\chi=\pi/4$ in the bottom panel of Figure 6.
@@
The PPA rotation over the full width at half maximum of the polarisation peak is evaluated for $\chi=\pi/4$ (the rotation is maximised for $\chi=\pi/4$). We obtain rotation values of $\Delta\phi$ = 3.5, 2.9, 2.6 degrees for $e=0.3$, $\Delta\phi$ = 16, 13, 12 degrees for $e=0.6$, corresponding to $\xi/\left\langle\theta_{c}/\theta_{j}\right\rangle=2,3,4$, respectively. While these values are evaluated for $\alpha=3$, the rotation remains insensitive to $\alpha$, with slightly larger values obtained for $\alpha=2$.
As discussed in Section 2, the polarisation degree curves exhibit a single peak even for top-hat, non-axisymmetric jets when the line of sight is close to the jet axis. In such cases, the position angle rotations are more pronounced, reaching at least, several tens degrees for our samples with $e=0.3,0.45$ or 0.6.
@@
Figure 7 illustrates polarisation signals when the line of sight passes through the jet core, corresponding to $\xi=0.5\theta_{c}/\theta_{j}$ and $\chi=\pi/4$. The polarisation curves resemble those of top-hat, non-axisymmetric jets (represented by the red solid line in the top panel of Figure 2), exhibiting two peaks without a complete polarisation dip in between. The trajectories in the $(q,u)$ space (the bottom panel) and the PPA rotation between two peaks show little sensitivity to $\alpha$. Specifically, for $\alpha=2$ and 3, we obtain $\Delta\phi\sim 79$ degrees and 80 degrees, respectively. In the limit $\alpha\gg 1$ (i.e., the top-hat model), the rotation angle increases slightly to $\Delta\phi\sim 83$ degrees. Overall, the top-hat model serve as a reasonable approximation of the polarisation signals when the line of sight intersects the jet core.

### v2

@@ 1 Introduction
Numerical simulations suggest such jet structures can arise from the interaction between the jet and the confining medium (Gottlieb et al., 2021). Specifically, this occurs as the jet breaks out through the stellar envelope in the collapsar model (e.g. Zhang et al. 2003; Lazzati & Begelman 2005; Gottlieb et al. 2021); or interacts with merger ejecta (e.g. Perego et al. 2014; Nativi et al. 2022). While most afterglow studies to date have focused on axisymmetric jets (either top-hat or structured), some have investigated azimuthal asymmetries in the energy and velocity distributions. An early example is Nakar & Oren 2004, who examined polarisation signals in the patchy-shell model, demonstrating that angular inhomogeneities in an expanding shock can lead to light-curve variability, accompanied by continuous, correlated changes in both the polarisation degree and position angle. More recently, Li et al. 2023 and Li et al. 2024, discuss azimuthal energy and velocity inhomogeneities produced by internal non-uniform magnetic dissipation processes or the precession of the central engine (Huang et al., 2019). Recent three-dimensional hydrodynamic simulations of jets in the aftermath of neutron star mergers (Gottlieb et al. 2022; Lamb et al. 2022) further highlight these complexities. These studies show that when a jet breaks out of the ejecta, the resulting outflow exhibits a power-law-like polar energy distribution with rotational inhomogeneity. Notably, the jet head is not perfectly circular, as often assumed in conventional theoretical models, instead, it tends to exhibit a more elliptical shape.
Non-axisymmetric or inhomogeneous structures have been primarily discussed in the context of prompt gamma-ray emission (e.g. Gill & Granot 2024; Lazar et al. 2009). This emission originates from the original ejecta from the central engine, which carries large-scale magnetic fields generated at the central source. In contrast, the magnetic fields in the afterglow-emitting blast wave are thought to originate from shock-driven instabilities, resulting in a highly tangled structure whose properties are largely governed by the anisotropy of the post-shock field. Detailed afterglow light-curve modelling can break degeneracies between jet structure, viewing geometry, and magnetic field anisotropy (Gill & Granot, 2020).
@@
We aim to investigate how jet axial asymmetry leaves an imprint on the temporal evolution of afterglow polarisation. To break axisymmetry, we model the global jet structure with an elliptical head. Jets emerging from a collapsing star are unlikely to naturally form an elliptical head, as this would require a non-axisymmetric stellar structure; in reality, stars are typically axisymmetric around their rotational axis. In double neutron-star mergers, however, the ejecta along the rotational axis originates from the neutron star collision and may display non-axisymmetric features. Nevertheless, it is worth exploring deviations from axisymmetry, since a perfectly circular jet head is itself an idealised case.
We do not claim that GRB jets generally have elliptical heads. Rather, we adopt this geometry as a toy model to explore departures from the standard top-hat jet. While there is no strong physical motivation for explicitly elliptical heads, the eccentricity provides a simple parameter to control the degree of asymmetry in the global jet structure. This serves as a straightforward starting point for investigating non-axisymmetric jets in this context.
We present our model for evaluating polarisation signals around a jet break in Section 2, where we illustrate the primary polarisation signature of a non-axisymmetric jet. We describe the distributions of polarisation signals when the line of sight of an observer intersects the jet head (or shock surface) at random locations in Section 3. In Section 4, we discuss the structured model in the context of non-axisymmetric jets. In Section 5 we present case studies. Finally, in Section 6, we give our conclusions and discussion.
@@ 2 Polarisation Behaviour at a Jet Break
Figure 2: Top panel: Polarisation as a function of time, with both quantities normalised to the the local polarisation degree $P_{0}$ and the characteristic jet-break time $t_{j}$
, for $e=0$ (blue dashed) and 0.6 (red solid). The first peaks, the troughs and second peaks in the polarisation light curves are marked by circles, crosses and squares, respectively. The line of sight is characterised by $\chi=1$ and $\xi=0.3$. Bottom panel: Same as the top panel, but showing the the temporal evolution of the polarisation position angle (PPA).
In the afterglow phase, magnetic fields in the blast wave are believed to be produced by shock instabilities, though the precise mechanisms responsible for their production in relativistic collisionless shocks remain poorly understood. Since the shock normal (aligned with the radial direction) represents a distinct orientation, the magnetic field components parallel and perpendicular to the shock normal could exhibit significantly different average strengths (Medvedev & Loeb 1999; Sari 1999; Ghisellini & Lazzati 1999). While many models assume that the post-shock field lies entirely within the shock plane, some observations indicate a more isotropic geometry (Gill & Granot 2020; Stringer & Lazzati 2020).
@@
where the local polarisation degree $P_{e}=P_{0}$ when the segment lies within the ring, and $P_{e}=0$ when the segment is inside the inner circle. The local luminosity, $dL$, is proportional to the area of the segment while also incorporating the constraints imposed by conditions (3), (5), and (6) in the previous paragraph. The angle $\phi_{e}$ denotes the angle between the positive x-axis (the semi-major axis) and the segment’s position vector, measured relative to the line of sight (Point A in Figure 1). Considering condition (6), $\Sigma dQ$, $\Sigma dU$ and $\Sigma dL$ are evaluated to estimate the net values of the parameters: $q=\Sigma dQ/\Sigma dL$ and $u=\Sigma dU/\Sigma dL$. The net polarisation degree $P$ and position angle $\phi$ are given by
@@
Figure 2 illustrates the differences in polarisation signals between axisymmetric jets and non-axisymmetric jets (in this paper, non-axisymmetric jets refer to those with elliptical-shaped heads). For axisymmetric jets, the polarisation degree curve (shown as the blue dashed line in the top panel) exhibits two peaks, with the polarisation dropping to zero between the peaks around the jet break time. As seen in the bottom panel, the polarisation position angle (PPA) $\phi$ rotates precisely by $\pi/2$ before and after the point where polarisation vanishes (Sari 1999; Ghisellini & Lazzati 1999). In contrast, jets with elliptical heads generally do not exhibit a complete polarisation drop between the two peaks ($P\neq 0$ at the trough). The PPA rotates more smoothly, though it still undergoes rapid changes near the trough (approximately at the jet break time). The change in PPA between the two peaks can significantly deviate from $\pi/2$; in this example, it is approximately $\Delta\phi=0.28\pi\sim 50.4$ degrees (it is more convenient to express angles in radians for theoretical discussion, while degrees are preferred in the context of observations. Unless otherwise specified, angles are given in radians).
@@ 3 The Characterisation of polarisation signals
Similarly, consider Point C, which is symmetric to Point A with respect to the x-axis. The fluid element distribution around Point C mirrors that of Point A but with a reflection of $\phi\leftrightarrow-\phi$ (flipping vertically with respect to the line passing through Point A and parallel to the x-axis). While the absolute values of the Stokes parameters at Point C match those at Point A, one of them flips sign, specifically (q,u)(q,-u). Nevertheless, the position angle rotation $|\Delta\phi-\pi/2|$ and the ratio $R_{pd}$ remain the same as at Point A.
To investigate intrinsically non-axisymmetric outflows, we model the asymmetric emitting surface with an elliptical cross section, recovering a conical jet when the eccentricity of the jet head is zero, while assuming that the surface dynamics (i.e., the radial motion of each fluid element) still follow the relativistic expansion described by the Blandford-McKee solution. As a result, the emitting surface possesses curvature. The ellipse represented in Figure 2 represents the projection of the jet head onto the sky. The opening angle of GRB jets or their core size is typically only several degrees. For such small angles, uniform sampling over projected area is approximately equivalent to uniform sampling over the solid angle of the emitting surface.
Figure 3 shows the distributions of $|\Delta\phi-\pi/2|$, derived from 5000 uniform sampled line of sight realisations across a quadrant of the elliptic jet head for eccentricity $e=0.3$, 0.45, and 0.6. For axisymmetric jets ($e=0$), the probability distribution is given by a delta function at 0, meaning the PPA rotation $\Delta\phi$ is always $\pi/2$. As eccentricity increases, the probability distribution (solid lines) peaks at larger deviation angles, with the distribution’s spread also increasing. The probability density distributions peak at $|\Delta\phi-\pi/2|$ = 4.3, 9.5 and 19 degrees, while their cumulative distributions reach 50% at 3.8, 8.9, 18 degrees, and 90% at 7.3, 17, 31 degrees for $e=0.3,0.45$ and 0.6, respectively. Since the semi-minor axis is smaller than the semi-major axis by a factor of $\sqrt{1-e^{2}}$, it is reduced by approximately 5%–20% (or equivalently by a factor of $\sim 0.8-0.95$) across this range of eccentricities. Our results show that even a mild deformation of the jet head leads to PPA rotations that deviate significantly from $\pi/2$.
@@
For even smaller offsets, one of the peaks, typically the first, becomes less pronounced. To determine whether a polarisation curve exhibits two peaks and to locate them, we have implemented a peak-finding algorithm similar to the one discussed in Li & Fenimore 1996 (see Appendix A for more details). For approximately, 4%, 7% and 11% of uniform random realisations for the $e=0.3$, 0.45 and 0.6, respectively, the polarisation curves exhibit only a single peak. These single-peak cases have been excluded from the statistical analysis of PPA rotation presented in Figs 3, 4 and 5.
Single-peak behaviour can be understood from the simple configuration where the line of sight passes through the jet centre (i.e., the origin; see Figure 1 for the following discussion). The visible region forms a ring centred at this point, with each fluid element emitting polarised light in the local radial direction. For a circular jet head ($e=0$), the system’s symmetry leads to complete cancellation of the polarisation signal. In contrast, for an eccentric jet head ($e\neq 0$), the initial net polarisation is zero, but as the jet decelerates, the top and bottom parts of the ring, emitting vertically polarised light, extend beyond the jet edge, breaking the symmetry. The net polarisation then becomes dominated by the left and right parts of the ring, whose emission is polarised in the horizontal direction. As deceleration proceeds, these segments continue to dominate, resulting in a constant PPA and a single peak in the polarisation curve. Although the probability of the line of sight passing exactly through the jet centre is zero, we find that the polarisation curve still exhibits a single peak, accompanied by a rotation of the PPA, when the line of sight lies close to the centre. As the viewing angle increases, a minor bump typically emerges during the rising phase and gradually evolves into a two-peak structure.
@@ 4 Non-axisymmetric structured jet
Figure 6: Polarisation signals for a line of sight intersecting the outer region, where the surface luminosity density declines following a power-law with an index of $\alpha=2$ (red) or 3 (blue). Two offset angles are considered: $\xi=2\theta_{c}/\theta_{j}$ (solid) and $4\theta_{c}/\theta_{j}$ (dashed). Top panel: temporal evolution of polarisation degree. Bottom panel: temporal evolution of polarisation position angle
. The model assumes $e=0.3$, $\chi=\pi/4$, and $\theta_{c}/\theta_{j}=0.13$.
Figure 7: Polarisation signals for a line of sight intersecting the jet core at $\xi=0.5\theta_{c}/\theta_{j}$. All jet properties are identical to those in Figure 6, except where specified. Top panel: polarisation degree curves. Bottom panel: Evolution of the polarisation position angle
. Curves are shown for $\alpha=2$ (red) and $\alpha=3$ (blue). The first peaks, the troughs and second peaks in the polarisation light curves are marked by circles, crosses and squares, respectively.
@@
In a structured jet, the jet break time for an off-axis observer is determined by the viewing angle $\theta_{view}$, rather than by the jet opening angle as in the top-hat jet model (Zhang & Mészáros 2002; Rossi et al. 2002). We estimate the characteristic jet-break time $t_{j}$ using a simple relation $\Gamma(t_{j})\sim 1/\theta_{eff}$, where $\theta_{eff}=\theta_{view}$ if the line of sight is in the outer region, and $\theta_{eff}=\theta_{c}$ if it is within the core.
@@
Figure 6 shows the polarisation signals when the line of sight falls in the outer region. As demonstrated by Rossi et al. 2004 and Birenbaum et al. 2024, the polarisation curves differ significantly from those of a top-hat jet, regardless of axial symmetry. This discrepancy arises from the non-uniform luminosity distribution in the visible region surrounding the line of sight (i.e., the ring), with the intense emission concentrated toward the jet axis. Consequently, the observed polarisation signal is primarily influenced by these bright segments, resulting in a single peak in the polarisation curve.
@@
The position angle evolution also differs from those of a top-hat jet. It consistently aligns with the direction toward the central jet axis (see assumption (4) in Section 2). When the line of sight falls on the jet head near its semi-major or semi-minor axis ($\chi\sim 0$ or $\pi/2$), the geometry becomes more symmetric, the position angle remains constant over time, consistent with the results for axisymmetric structured jets (Rossi et al., 2004). However, for a general azimuthal angle $\chi$, the eccentricity of the brightness contour lines leads to a slight evolution of the position angle around the peak of the polarisation degree. This effect is illustrated for $\chi=\pi/4$ in the bottom panel of Figure 6.
@@
The PPA rotation over the full width at half maximum of the polarisation peak is evaluated for $\chi=\pi/4$ (the rotation is maximised for $\chi=\pi/4$). We obtain rotation values of $\Delta\phi$ = 3.5, 2.9, 2.6 degrees for $e=0.3$, $\Delta\phi$ = 16, 13, 12 degrees for $e=0.6$, corresponding to $\xi/\left\langle\theta_{c}/\theta_{j}\right\rangle=2,3,4$, respectively. While these values are evaluated for $\alpha=3$, the rotation remains insensitive to $\alpha$.
As discussed in Section 3, the polarisation degree curve can exhibit a single peak even for top-hat, non-axisymmetric jets when the line of sight is close to the jet axis. In such cases, the PPA typically rotates between a minor bump and the main peak, reaching a local maximum or minimum at the peak of the polarisation curve. The rotation rate of the PPA often differs markedly between the rising and decaying phases. In contrast, for structured jets, the PPA evolves more smoothly and continuously across the peak.
@@
Figure 7 illustrates polarisation signals when the line of sight passes through the jet core, corresponding to $\xi=0.5(\theta_{c}/\theta_{j})$ and $\chi=\pi/4$. The polarisation curves resemble those of top-hat, non-axisymmetric jets (represented by the red solid line in the top panel of Figure 2), exhibiting two peaks without a complete polarisation dip in between. The temporal evolution of the PPA (shown in the bottom panel) is largely insensitive to the value of $\alpha$. When $\alpha=2$ and 3, we obtain the PPA rotation between the two peaks is $\Delta\phi\sim 79$ degrees and 80 degrees, respectively. In the limit $\alpha\gg 1$ (i.e., the top-hat model), the rotation angle increases slightly to $\Delta\phi\sim 83$ degrees. Overall, the top-hat model serve as a reasonable approximation of the polarisation signals when the line of sight intersects the jet core.